OpenStudy (anonymous):

lim x->0 12(1-cosx)/x^2 what is the limit?

OpenStudy (turingtest):

try multiplying top and bottom by the conjugate of 1-cosx

OpenStudy (anonymous):

This might involve using the known limit, \(\displaystyle\lim_{x\to0}\frac{1-\cos x}{x}=0\).

OpenStudy (turingtest):

I wasn't able to work that into it. Multiply top and bottom by 1+cosx, trig identity, then lim x->0 sinx/x is all you need